Solve each equation. Check for extraneous solutions.

1. [7x+2]-2=7x
2.(-2-x)^0.5-x=2

Answer by Alan3354(21605)

Solve each equation. Check for extraneous solutions.
1. [7x+2]-2=7x
7x = 7x
There's no unique solution, x can be any value.
-----------------------------
2.(-2-x)^0.5-x=2
Square both sides
-2-x = 4
x = -6